# Local certification of geometric graph classes

### v4

@@ 7.5. Tied-arch bridges
In this section, $x>0$ and $0\leq\delta\leq 1$ are real numbers, and whenever we use the $O(\cdot)$ notation, we implicitly assume that $x\to\infty$ (see for instance the terms $O(1/x)$ and $O(\sqrt{\delta/x})$ in the statement of the next lemma). Let $abc$ be a triangle such that $d_{2}(a,b)=x$ and $d_{2}(a,c)+d_{2}(b,c)=x+\delta$, with $|d_{2}(a,c)-d_{2}(b,c)|\leq 1$. Assume by symmetry that $d_{2}(a,c)\leq d_{2}(b,c)$. In particular $\tfrac{1}{2}(x+\delta-1)\leq d_{2}(a,c)\leq\tfrac{1}{2}(x+\delta)$ and $\tfrac{1}{2}(x+\delta)\leq d_{2}(b,c)\leq\tfrac{1}{2}(x+\delta+1)$. Let $c^{\prime}$ be a point such that $d_{2}(a,c^{\prime})\leq d_{2}(a,c)$, $d_{2}(b,c^{\prime})\leq d_{2}(b,c)$, and $d_{2}(c,c^{\prime})\leq\rho$, for some constant $\rho=O(1)$. See Figure 14 for an illustration.
@@
Consider a graph $G_{0}$ given by Observation 7.5 with parameter $2k$. The unit-disk graph $G_{0}$ thus contains $O(k^{1+\varepsilon})$ vertices and has two vertices $u^{\prime}$ and $v^{\prime}$ such that in any unit-disk embedding $12k\leq d_{2}(u^{\prime},v^{\prime})\leq 2k+$ 1 and $u^{\prime}$ and $v^{\prime}$ are joined by a path $P^{\prime}$ of length less than $12k+$ 1.
@@
We consider a vertex $w$ of $P$, which divides $P$ into two consecutive paths $P_{0}$ and $P_{1}$, whose lengths differ by at most 1. We consider a unit-disk embedding of $G_{0}$ and look at the perpendicular bisector of the line-segment $[u,v]$. By connectivity, this line intersects $G_{0}-P$. Let $z$ be the first vertex of $G_{0}-P$ whose unit-disk is intersected by this line (if several such vertices exist we pick one arbitrarily); by the properties above, $z$ lies at distance at least $\Omega(k)$ from $w$. As in the construction of quasi-rigid graphs above Observation 7.5, we consider a point set corresponding to a unit-disk embedding of $G_{0}$ where the Euclidean distance $m^{*}$ between $z$ and $w$ is minimized and add, along the segment $[zw]$, $\lceil m^{*}\rceil-1$ evenly spaced new points. Observe that in the resulting unit-disk graph $G_{1}$ the newly added vertices correspond to a path $Q$ of minimum length between $z$ and $w$ (and possibly some extra edges between this path and $V(G_{0})$). Therefore we have $m^{*}\leq d_{2}(z,w)\leq\lceil m^{*}\rceil$ in any unit-disk embedding of $G_{1}$. We iterate this procedure recursively on $P_{0}$ and $P_{1}$, creating 4 consecutive subpaths $P_{00},P_{01},P_{10},P_{11}$ of $P$, and two new paths $Q_{0}$ and $Q_{1}$ joining the new midpoints to $G-P$. More precisely, for any $i\geq 0$, consider a unit-disk embedding of $G_{i}$, and any subpath $P_{\mathbf{x}}$ of $P$ with $\mathbf{x}\in\{0,1\}^{i}$ in $G_{i}$. Let $a,b$ denote the endpoints of $P_{\mathbf{x}}$. Then in $G_{i+1}$, $P_{\mathbf{x}}$ is split between $P_{\mathbf{x}0}$ (with endpoints $a$ and $c^{\prime}$) and $P_{\mathbf{x}1}$ (with endpoints $c^{\prime}$ and $b$), where $c^{\prime}$ is a midpoint of $P_{\mathbf{x}}$, which is then joined (in the way described above) via some path $Q_{\mathbf{x}}$ of minimal length to some vertex $z$ of $G_{i}-P$ lying at distance at most $\tfrac{1}{2}$ from the perpendicular bisector of the line segment $[ab]$. We precise that at each recursive step $\mathbf{x}\in\{0,1\}^{i}$ , the paths $Q_{\mathbf{x}0}$ and $Q_{\mathbf{x}1}$ are added one after the other: we first add a path $Q_{\mathbf{x}0}$ of minimal length (such that the resulting graph is still a unit-disk graph), and then we add a path $Q_{\mathbf{x}1}$ of minimum length (such that the resulting graph is still a unit-disk graph). Note that by construction $G_{i+1}$ is a unit-disk graph. See Figure 15 for a picture of $G_{3}$.
@@
Consider a unit-disk embedding of a unit-disk graph $G$. For $\ell\geq 2$ and $\gamma\geq 0$, we say that a path $P=v_{1},\ldots,v_{\ell+1}$ of length $\ell$ in $G$ is $\gamma$-regular if the following holds: If we place $\ell+1$ evenly spaced points $u_{1},\ldots,u_{\ell+1}$ on the line-segment $[v_{1},v_{\ell+1}]$ with $u_{1}=v_{1}$ and $u_{\ell+1}=v_{\ell+1}$, then for any $11\leq i\leq\ell+$ 1, $d_{2}(u_{i},v_{i})\leq\gamma$. For $s\leq\ell+1$, we say that $P$ is $(\gamma,s)$-regular if the above holds for any $1\leq i\leq s$ (instead of $11\leq i\leq\ell+$ 1, so that $(\gamma,\ell+1)$-regular is the same as $\gamma$-regular for a path of length $\ell$) . In other words, this means that the first $s$ vertices of $P$ are close to their ideal location on the segment connecting the two endpoints of $P$.
@@
By Lemma 7.8, there is $\mathbf{x}\in\{0,1\}^{t}$ such that the length of the path $P_{\mathbf{x}}$ in $G_{t}$ with endpoints $a$ and $b$ differs from $d_{2}(a,b)$ by $\delta:=O(2^{t}\cdot k^{2^{-t}-1})$. As $P_{\mathbf{x}}$ has length $\Theta(2^{-t}k)$, it follows from Lemma 7.9 that $P_{\mathbf{x}}$ is $\gamma$-regular with $\gamma=O\left(\sqrt{k^{2^{-t}}}\right)=O\left(k^{2^{-t-1}}\right)$. As $k^{2^{-t}}\leq 2^{\log k\cdot\log^{-1}k}\leq 2$, $\delta=O(k^{-1}\log k)$ and $\gamma=O(1)$, as desired. ∎
@@ 7.6. Corridors
A unit-disk embedding of an $r$-corridor is depicted in Figure 16 (bottom). The second property follows from the fact that there is a unique planar graph embedding (up to reflection, which follows from Theorem 2.2), and the 4 neighbors of $u$ (and $v$) are non-adjacent, and thus the minimum angle between $ux_{1}$ and $uy_{1}$ is at least $\pi$ (and similarly for the angle between $vx_{r+1}$ and $vy_{r+1}$). ∎
@@
Consider a unit-disk graph $G$ embedded in the plane with a $\delta$-tight path $P^{\prime}$, for $\delta\leq 1$. Assume that $P^{\prime}$ has a subpath $P$ of length $|P|\geq|P^{\prime}|/2$, with endpoints $u,v$, such that in any unit-disk embedding of $G$, $P$ lies at Euclidean distance at least $4$ from $G-P^{\prime}$. As a subpath of $P^{\prime}$, $P$ is also $\delta$-tight. Let $G^{\prime}$ be the graph obtained from $G$ by deleting the internal vertices of $P$ and adding a copy of a $|P|$-corridor $C$ with ends $u$ and $v$. We say that we have replaced the path $P$ by the corridor $C$ in $G$.
@@ 7.7. Decorating the corridors
This follows from the definition of a corridor: the purpose of the vertices $z_{1}$ and $z_{r+1}$ is to make sure that the line segments $[x_{1},x_{r+1}]$ and $[y_{1},y_{r+1}]$ are at Euclidean distance at least $\sqrt{3}$ from each other, which allows one of the two 3-vertex paths to be placed inside the corridor while the other is placed outside, for each location $i$, as illustrated in Figure 16 (bottom). ∎
@@ 7.8. Disjointness-expressivity of unit-disk graphs
By construction, $G^{\prime}_{A,B}$ is the graph $g(L(A),R(B))$ obtained from $L(A)$ and $R(B)$ by gluing them along $S$. As there are $\log k=O(\log N)$ corridors and the closed neighborhood of each vertex of $S$ contains 4 vertices, the size of $N[S]$ is $O(\log N)$. The fact that $G^{\prime}_{A,B}$ is a unit-disk graph if and only if $A$ and $B$ are disjoint is a direct consequence of Lemma 7.15. ∎
@@ Acknowledgement
Consider a $(k,\delta)$-almost-equilateral embedding of a stripe $S_{\ell}$ with $k\gg\delta$. Up to exchanging the roles of $u_{i}$ and $v_{i}$, the Euclidean distance between $u_{0}$ and $u_{\ell}$ is minimized when the distances $d_{2}(u_{i},u_{i+1})$ ($10\leq i\leq\ell-$ 1) are minimized and the angles $\angle u_{i-1}u_{i}u_{i+1}$ ($11\leq i\leq\ell-$ 1) are maximized. Since all the constraints on these distances and angles are local, in an embedding minimizing $d_{2}(u_{0},v_{\ell})$ all these lengths are equal, and all these angles are equal. It follows that the Menger curvature of all consecutive triples is the same (and only depends on $k$ and $\delta$, since the local constraints only depend on $k$ and $\delta$). ∎
@@
Consider a $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ where $d_{2}(u_{0},v_{\ell})\leq m+1$. The second part of the statement simply follows from the fact that in any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$, the diameter of the corresponding point set is at most $(k+\delta)\ell$, so we only need to prove the first part of the statement. We only give the main steps of the proof and omit the technical computations. By Lemma 7.3, for each $0\leq i<j\leq\ell$, the Euclidean distance between $u_{i}$ and $u_{j}$ is at least the distance between the endpoints of some circular arc of radius $\rho$ that subtends an angle $\tfrac{j-i}{\ell}\cdot\pi$, that is $d_{2}(u_{i},u_{j})\geq 2\rho\sin(\tfrac{j-i}{\ell}\cdot\tfrac{\pi}{2})$. It follows that for any $0\leq i\leq\ell$, $d_{2}(u_{0},u_{i})\geq 2\rho\sin(\alpha)$ and $d_{2}(u_{0},u_{i})\geq 2\rho\cos(\alpha)$, where $\alpha:=\tfrac{i\pi}{2\ell}$. We can use this to estimate the area of the triangle $u_{0}u_{i}u_{\ell}$
  1. (1)
    when $u_{0},u_{i},u_{\ell}$ all lie on a circle of radius $\rho$ and $d_{2}(u_{0},u_{\ell})=2\rho$, and
  2. (2)
    when we only require that $d_{2}(u_{0},u_{\ell})=2\rho+O(k)$.
Comparing the two areas (more precisely, in the second case, we only obtain a lower bound on the area), we can estimate the distance between $u_{i}$ and the segment $[u_{0}u_{\ell}]$ in both cases and then show that these distances differ by at most $O(k)$. We use this to conclude that in the second case, $u_{i}$ lies at distance $O(k)$ from the circle of radius $\rho$ centered in $c$. ∎
@@
Consider a unit-disk graph embedding of $T_{n}$, and let $P$ be the $(8n+16)$-gon corresponding to the vertices of $C$. We denote by $P_{1}$ and $P_{2}$ the two polygonal subchains of $P$ with endpoints $c$ and $c^{\prime}$. Let $R_{1}$ be the region bounded by $P_{1}$ and the line-segment $[c,c^{\prime}]$, and let $R_{2}$ be the region bounded by $P_{2}$ and the line-segment $[c,c^{\prime}]$. It might be the case that one of the two regions contains the other, but in any case the region $R$ bounded by $P$ is contained in the union of $R_{1}$ and $R_{2}$.
Note that $T_{n}$ has an independent set $S$ of size at least $(2n+2)^{2}$ (the set of red vertices, or the set of blue vertices in Figure 12). Discard the vertices of $S$ that are at Euclidean distance at most $\tfrac{1}{2}$ from $P$ or from the line-segment $[c,c^{\prime}]$ (there are at most $O(n)$ such vertices), and let $S^{\prime}$ be the resulting independent set (of size $(2n+2)^{2}-O(n)$). The disks of radius $\tfrac{1}{2}$ centered in the points corresponding to $S^{\prime}$ are pairwise disjoint, and all contained in $R$, and thus contained in $R_{1}$ or $R_{2}$. It follows that one of $R_{1}$ or $R_{2}$ (say $R_{1}$), contains at least $\tfrac{1}{2}(2n+2)^{2}-O(n)=2(n+1)^{2}-O(n)$ disjoint disks of radius $\tfrac{1}{2}$, and thus has area $A(R_{1})\geq\tfrac{\pi}{4}(2(n+1)^{2}-O(n))=\tfrac{\pi}{2}(n+1)^{2}-O(n)$. Let $d:=d_{2}(c,c^{\prime})$. Note that $R_{1}$ and $R_{2}$ have perimeter at most $4n+8+d$, and thus by the isoperimetric inequality in the plane, $A(R_{1})\leq\tfrac{1}{4\pi}(4n+8+d)^{2}$. It follows that $(4n+d+8)^{2}\geq 2\pi^{2}(n+1)^{2}-O(n)$, and thus $d\geq(\pi\sqrt{2}-4)n-O(\sqrt{n})=\Omega(n)$, as desired. ∎
@@
Let $y$ be the length of the altitude of $abc$ passing through $c$. Note that $y$ is maximized when $d_{2}(a,c)=d_{2}(b,c)$, and thus by Pythagoras’ theorem, $y\leq\sqrt{\delta x/2}+O(x^{-1/2})$. Note that subject to the conditions above, $\min\{d_{2}(a,c)-d_{2}(a,c^{\prime}),d_{2}(b,c)-d_{2}(b,c^{\prime})\}$ is maximized when $d_{2}(c,c^{\prime})=\rho$. Let $c^{*}$ be the point of the altitude of $abc$ passing through $c$ such that $d_{2}(c,c^{*})=\rho$. This altitude divides $abc$ into two triangles, say $T_{a}$ containing $a$ and $T_{b}$ containing $b$. If $c^{\prime}\in T_{a}$ then $d_{2}(b,c^{\prime})\geq d_{2}(b,c^{*})$, and if $c^{\prime}\in T_{b}$ then $d_{2}(a,c^{\prime})\geq d_{2}(a,c^{*})$. Thus it suffices to prove that $d_{2}(a,c)-d_{2}(a,c^{*})\leq\rho\sqrt{2\delta/x}+O(1/x)$ and $d_{2}(b,c)-d_{2}(b,c^{*})\leq\rho\sqrt{2\delta/x}+O(1/x)$. The two cases being symmetric, we only prove the former. Let $z=d_{2}(a,c)$. By Pythagoras’ theorem,
@@
Let $x:=d_{2}(a,b)$, and let $c\in\mathbb{R}^{2}$ be such that $d_{2}(a,c)=|P_{\mathbf{x}0}|$ and $d_{2}(b,c)=|P_{\mathbf{x}1}|$. There are two choices for $c$ which are symmetric with respect to the line-segment $[a,b]$, and we take the one which is closest from $z$. The path $P_{\mathbf{x}}$ is $\delta$-tight so $d_{2}(a,c)+d_{2}(c,b)=x+\delta$. As $c^{\prime}$ is a midpoint of $P_{\mathbf{x}}$, $|d_{2}(a,c)-d_{2}(b,c)|\leq 1$. Hence, in order to apply Lemma 7.6, which directly gives us the desired result, we only need to prove that $\rho:=d_{2}(c,c^{\prime})=O(1)$.

### v5

@@ 7.5. Tied-arch bridges
In this section, $x>0$ and $0\leq\delta\leq 1$ are real numbers, and whenever we use the $O(\cdot)$ notation, we implicitly assume that $x\to\infty$ (see for instance the terms $O(1/x)$ and $O(\sqrt{\delta/x})$ in the statement of the next lemma). Let abcabc be a triangle such that $d_{2}(a,b)=x$ and $d_{2}(a,c)+d_{2}(b,c)=x+\delta$, with $|d_{2}(a,c)-d_{2}(b,c)|\leq 1$. Assume by symmetry that $d_{2}(a,c)\leq d_{2}(b,c)$. In particular $\tfrac{1}{2}(x+\delta-1)\leq d_{2}(a,c)\leq\tfrac{1}{2}(x+\delta)$ and $\tfrac{1}{2}(x+\delta)\leq d_{2}(b,c)\leq\tfrac{1}{2}(x+\delta+1)$. Let $c^{\prime}$ be a point such that $d_{2}(a,c^{\prime})\leq d_{2}(a,c)$, $d_{2}(b,c^{\prime})\leq d_{2}(b,c)$, and $d_{2}(c,c^{\prime})\leq\rho$, for some constant $\rho=O(1)$. See Figure 14 for an illustration.
@@
Consider a graph $G_{0}$ given by Observation 7.5 with parameter $2k$. The unit-disk graph $G_{0}$ thus contains $O(k^{1+\varepsilon})$ vertices and has two vertices $u^{\prime}$ and $v^{\prime}$ such that in any unit-disk embedding $2k\leq d_{2}(u^{\prime},v^{\prime})\leq 2k+1$ and $u^{\prime}$ and $v^{\prime}$ are joined by a path $P^{\prime}$ of length less than $2k+1$.
@@
It is helpful to have a quick look at Figure 15 before reading this section. We consider a vertex $w$ of $P$, which divides $P$ into two consecutive paths $P_{0}$ and $P_{1}$, whose lengths differ by at most 1. We consider a unit-disk embedding of $G_{0}$ and look at the perpendicular bisector of the line-segment $[u,v]$. By connectivity, this line intersects $G_{0}-P$. Let $z$ be the first vertex of $G_{0}-P$ whose unit-disk is intersected by this line (if several such vertices exist we pick one arbitrarily); by the properties above, $z$ lies at distance at least $\Omega(k)$ from $w$. As in the construction of quasi-rigid graphs above Observation 7.5, we consider a point set corresponding to a unit-disk embedding of $G_{0}$ where the Euclidean distance $m^{*}$ between $z$ and $w$ is minimized and add, along the segment $[zw]$, $\lceil m^{*}\rceil-1$ evenly spaced new points. Observe that in the resulting unit-disk graph $G_{1}$ the newly added vertices correspond to a path $Q$ of minimum length between $z$ and $w$ (and possibly some extra edges between this path and $V(G_{0})$). Therefore we have $m^{*}\leq d_{2}(z,w)\leq\lceil m^{*}\rceil$ in any unit-disk embedding of $G_{1}$. We iterate this procedure recursively on $P_{0}$ and $P_{1}$, creating 4 consecutive subpaths $P_{00},P_{01},P_{10},P_{11}$ of $P$, and two new paths $Q_{0}$ and $Q_{1}$ joining the new midpoints to $G_{0}-P$. More precisely, for any $i\geq 0$, consider a unit-disk embedding of $G_{i}$, and any subpath $P_{\mathbf{x}}$ of $P$ with $\mathbf{x}\in\{0,1\}^{i}$ in $G_{i}$. Let $a,b$ denote the endpoints of $P_{\mathbf{x}}$. Then in $G_{i+1}$, $P_{\mathbf{x}}$ is split between $P_{\mathbf{x}0}$ (with endpoints $a$ and $c^{\prime}$) and $P_{\mathbf{x}1}$ (with endpoints $c^{\prime}$ and $b$), where $c^{\prime}$ is a midpoint of $P_{\mathbf{x}}$, which is then joined (in the way described above) via some path $Q_{\mathbf{x}}$ of minimal length to some vertex $z_{\mathbf{x}}$ of $G_{i}-P$ lying at distance at most $\tfrac{1}{2}$ from the perpendicular bisector of the line segment $[ab]$. We precise that at each recursive step $\mathbf{x}\in\{0,1\}^{i}$ , the paths $Q_{\mathbf{x}0}$ and $Q_{\mathbf{x}1}$ are added one after the other: we first add a path $Q_{\mathbf{x}0}$ of minimal length (such that the resulting graph is still a unit-disk graph), and then we add a path $Q_{\mathbf{x}1}$ of minimum length (such that the resulting graph is still a unit-disk graph). Note that by construction $G_{i+1}$ is a unit-disk graph. See Figure 15 for a picture of $G_{3}$.
@@
Consider a unit-disk embedding of a unit-disk graph $G$. For $\ell\geq 2$ and $\gamma\geq 0$, we say that a path $P=v_{1},\ldots,v_{\ell+1}$ of length $\ell$ in $G$ is $\gamma$-regular if the following holds: If we place $\ell+1$ evenly spaced points $u_{1},\ldots,u_{\ell+1}$ on the line-segment $v_{1},v_{\ell+1}]$ with $u_{1}=v_{1}$ and $u_{\ell+1}=v_{\ell+1}$, then for any $1\leq i\leq\ell+1$, $d_{2}(u_{i},v_{i})\leq\gamma$. For $s\leq\ell+1$, we say that $P$ is $(\gamma,s)$-regular if the above holds for any $1\leq i\leq s$ (instead of $1\leq i\leq\ell+1$, so that $(\gamma,\ell+1)$-regular is the same as $\gamma$-regular for a path of length $\ell$) . In other words, this means that the first $s$ vertices of $P$ are close to their ideal location on the segment connecting the two endpoints of $P$.
@@
By Lemma 7.8, there is $\mathbf{x}\in\{0,1\}^{t}$ such that the length of the path $P_{\mathbf{x}}$ in $G_{t}$ with endpoints $a$ and $b$ differs from $d_{2}(a,b)$ by $\delta:=O(2^{t}\cdot k^{2^{-t}-1})$. As $P_{\mathbf{x}}$ has length $\Theta(2^{-t}k)$, it follows from Lemma 7.9 that $P_{\mathbf{x}}$ is $\gamma$-regular with $\gamma=O\left(\sqrt{k^{2^{-t}}}\right)=O\left(k^{2^{-t-1}}\right)$. As $k^{2^{-t}}\leq 2^{\log k\cdot\log^{-1}k}=2$, $\delta=O(k^{-1}\log k)$ and $\gamma=O(1)$, as desired. ∎
@@ 7.6. Corridors
A unit-disk embedding of an $r$-corridor is depicted in Figure 16 (bottom). The second property follows from the fact that there is a unique planar graph embedding (up to reflection, which follows from Theorem 2.2), and the 4 neighbors of $u$ (and $v$) are non-adjacent, and thus the angle between $ux_{1}$ and $uy_{1}$ is at least $\pi$ (and similarly for the angle between $vx_{r+1}$ and $vy_{r+1}$). ∎
@@
Consider a unit-disk graph $G$ embedded in the plane with a $\delta$-tight path $P^{\prime}$, for $\delta\leq 1$. Assume that $P^{\prime}$ has a subpath $P$ of length $|P|\geq|P^{\prime}|/2$, with endpoints $u,v$, such that in any unit-disk embedding of $G$, $P$ lies at Euclidean distance at least $4$ from $G-P^{\prime}$. As a subpath of $^{\prime}, PP$ is also $\delta$-tight. Let $G^{\prime}$ be the graph obtained from $G$ by deleting the internal vertices of $P$ and adding a copy of a $|P|$-corridor $C$ with ends $u$ and $v$. We say that we have replaced the path $P$ by the corridor $C$ in $G$.
@@ 7.7. Decorating the corridors
This follows from the definition of a corridor: the purpose of the vertices $z_{1}$ and $z_{r+1}$ is to make sure that the line segments $x_{1},x_{r+1}]$ and $y_{1},y_{r+1}]$ are at Euclidean distance at least $\sqrt{3}$ from each other, which allows one of the two 3-vertex paths to be placed inside the corridor while the other is placed outside, for each location $i$, as illustrated in Figure 16 (bottom). ∎
@@ 7.8. Disjointness-expressivity of unit-disk graphs
By construction, $G^{\prime}_{A,B}$ is the graph $g(L(A),R(B))$ obtained from $L(A)$ and $R(B)$ by gluing them along $S$. As there are $\log k=O(\log N)$ corridors and the closed neighborhood of each vertex of $S$ contains 4 vertices, the size of $N[S]$ is $O(\log N)$. Moreover, the subgraph induced by the closed neighborhood of each vertex of $S$ is independent of $A$ and $B$, and as the vertices of $S$ are pairwise at distance at least 4 apart, the subgraph induced by the closed neighborhood of $S$ is the disjoint union of all subgraphs induced by $N[s]$, for $s\in S$, which is independent of $A$ and $B$. The fact that $G^{\prime}_{A,B}$ is a unit-disk graph if and only if $A$ and $B$ are disjoint is a direct consequence of Lemma 7.15. ∎
@@ 8. Open problems
As a final remark, it was suggested by one of the reviewers that the lower bound techniques used in the paper can be used to prove lower bounds for the recognition problem in the studied classes in different models of computation, namely in the CONGEST model of distributed computing where messages exchanged by nodes have logarithmic size, and in the streaming model.
@@ Acknowledgement
Consider a $(k,\delta)$-almost-equilateral embedding of a stripe $S_{\ell}$ with $k\gg\delta$. Up to exchanging the roles of $u_{i}$ and $v_{i}$, the Euclidean distance between $u_{0}$ and $u_{\ell}$ is minimized when the distances $d_{2}(u_{i},u_{i+1})$ ($0\leq i\leq\ell-1$) are minimized and the angles $\angle u_{i-1}u_{i}u_{i+1}$ ($1\leq i\leq\ell-1$) are maximized. Since all the constraints on these distances and angles are local, in an embedding minimizing $d_{2}(u_{0},v_{\ell})$ all these lengths are equal, and all these angles are equal. It follows that the Menger curvature of all consecutive triples is the same (and only depends on $k$ and $\delta$, since the local constraints only depend on $k$ and $\delta$). ∎
@@
Consider a $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ where $d_{2}(u_{0},v_{\ell})\leq m+1$. The second part of the statement simply follows from the fact that in any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$, the diameter of the corresponding point set is at most $(k+\delta)\ell$, so we only need to prove the first part of the statement. By Lemma 7.2, for each $0\leq i<j\leq\ell$, the Euclidean distance between $u_{i}$ and $u_{j}$ is at least the distance between the endpoints of some circular arc of radius $\rho$ that subtends an angle $\tfrac{j-i}{\ell}\cdot\pi$, that is $d_{2}(u_{i},u_{j})\geq 2\rho\sin(\tfrac{j-i}{\ell}\cdot\tfrac{\pi}{2})$. It follows that for any $0\leq i\leq\ell$, $d_{2}(u_{0},u_{i})\geq 2\rho\sin(\alpha)$ and $d_{2}(u_{i},u_{\ell})\geq 2\rho\cos(\alpha)$, where $\alpha:=\tfrac{i\pi}{2\ell}$. On the other hand, note that $d_{2}(u_{0},v_{\ell})\leq m+1\leq 2\rho+O(k)$, where the second inequality follows from the definition of $\ell$. By Apollonius’ theorem,
$d_{2}(u_{i},c)^{2}$ $=$ $\tfrac{1}{2}(d_{2}(u_{0},u_{i})^{2}+d_{2}(u_{i},v_{\ell})^{2})-d_{2}(u_{0},c)^{2}$
$\geq$ $2\rho^{2}\sin^{2}\alpha+2\rho^{2}\cos^{2}\alpha-2k\rho\cos\alpha+2k^{2}-(\rho^{2}+O(\rho k)+O(k^{2}))$
$\geq$ $\rho^{2}-2k\rho-O(k^{2})$
$\geq$ $\rho^{2}(1-O(k/\rho+k^{2}/\rho^{2}))$
Figure 17. Illustration of the proof of Lemma 7.3.
It follows that
$d_{2}(u_{i},c)\geq\rho\cdot(1-O(k/\rho+k^{2}/\rho^{2}))\geq\rho-O(k+k^{2}/\rho)\geq\rho-O(k),$
where the final inequality follows from the fact that $k^{2}/\rho=O(k^{2}/m)=O(\delta)=o(k)$. A similar computation applies to the points $v_{i}$, which concludes the proof. ∎
@@
Consider a unit-disk graph embedding of $T_{n}$, and let $P$ be the $(8n+16)$-gon corresponding to the vertices of $C$. We denote by $P_{1}$ and $P_{2}$ the two polygonal subchains of $P$ with endpoints $c$ and $c^{\prime}$. Let $R_{1}$ be the region bounded by $P_{1}$ and the line-segment $c,c^{\prime}]$, and let $R_{2}$ be the region bounded by $P_{2}$ and the line-segment $c,c^{\prime}]$. It might be the case that one of the two regions contains the other, but in any case the region $R$ bounded by $P$ is contained in the union of $R_{1}$ and $R_{2}$.
Note that $T_{n}$ has an independent set $S$ of size at least $(2n+2)^{2}$ (the set of red vertices, or the set of blue vertices in Figure 12). Discard the vertices of $S$ that are at Euclidean distance at most $\tfrac{1}{2}$ from $P$ or from the line-segment $c,c^{\prime}]$ (there are at most $O(n)$ such vertices), and let $S^{\prime}$ be the resulting independent set (of size $(2n+2)^{2}-O(n)$). The disks of radius $\tfrac{1}{2}$ centered in the points corresponding to $S^{\prime}$ are pairwise disjoint, and all contained in $R$, and thus contained in $R_{1}$ or $R_{2}$. It follows that one of $R_{1}$ or $R_{2}$ (say $R_{1}$), contains at least $\tfrac{1}{2}(2n+2)^{2}-O(n)=2(n+1)^{2}-O(n)$ disjoint disks of radius $\tfrac{1}{2}$, and thus has area $A(R_{1})\geq\tfrac{\pi}{4}(2(n+1)^{2}-O(n))=\tfrac{\pi}{2}(n+1)^{2}-O(n)$. Let $d:=d_{2}(c,c^{\prime})$. Note that $R_{1}$ and $R_{2}$ have perimeter at most $4n+8+d$, and thus by the isoperimetric inequality in the plane, $A(R_{1})\leq\tfrac{1}{4\pi}(4n+8+d)^{2}$. It follows that $(4n+d+8)^{2}\geq 2\pi^{2}(n+1)^{2}-O(n)$, and thus $d\geq(\pi\sqrt{2}-4)n-O(\sqrt{n})=\Omega(n)$, as desired. ∎
@@
Let $y$ be the length of the altitude of abcabc passing through $c$. Note that $y$ is maximized when $d_{2}(a,c)=d_{2}(b,c)$, and thus by Pythagoras’ theorem, $y\leq\sqrt{\delta x/2}+O(x^{-1/2})$. Note that subject to the conditions above, $\min\{d_{2}(a,c)-d_{2}(a,c^{\prime}),d_{2}(b,c)-d_{2}(b,c^{\prime})\}$ is maximized when $d_{2}(c,c^{\prime})=\rho$.
Figure 18. Illustration of the proof of Lemma 7.6.
Let $c^{*}$ be the point of the altitude of abcabc passing through $c$ such that $d_{2}(c,c^{*})=\rho$. See Figure 18 for an illustration. This altitude divides abcabc into two triangles, say $T_{a}$ containing $a$ and $T_{b}$ containing $b$. If $c^{\prime}\in T_{a}$ then $d_{2}(b,c^{\prime})\geq d_{2}(b,c^{*})$, and if $c^{\prime}\in T_{b}$ then $d_{2}(a,c^{\prime})\geq d_{2}(a,c^{*})$. Thus it suffices to prove that $d_{2}(a,c)-d_{2}(a,c^{*})\leq\rho\sqrt{2\delta/x}+O(1/x)$ and $d_{2}(b,c)-d_{2}(b,c^{*})\leq\rho\sqrt{2\delta/x}+O(1/x)$. The two cases being symmetric, we only prove the former. Let $z=d_{2}(a,c)$. By Pythagoras’ theorem,
@@
Let $x:=d_{2}(a,b)$, and let $c\in\mathbb{R}^{2}$ be such that $d_{2}(a,c)=|P_{\mathbf{x}0}|$ and $d_{2}(b,c)=|P_{\mathbf{x}1}|$. This defines (at most) 2 points which are symmetric with respect to the line-segment $[a,b]$, and we define $c$ as the one which is closest from $z_{\mathbf{x}}$. See Figure 19 for an illustration. The path $P_{\mathbf{x}}$ is $\delta$-tight so $d_{2}(a,c)+d_{2}(c,b)=x+\delta$. As $c^{\prime}$ is a midpoint of $P_{\mathbf{x}}$, $|d_{2}(a,c)-d_{2}(b,c)|\leq 1$. Hence, in order to apply Lemma 7.6, which directly gives us the desired result, we only need to prove that $\rho:=d_{2}(c,c^{\prime})=O(1)$.
Figure 19. Illustration of the proof of Corollary $7.$